OpenStudy (anonymous):

in 2000 the population of a state was 4 million and growing continuously at a rate of 3.3% per year. a). write a function that models the state's population in millions x years after 2000 b). graph f in[0,10] by [4,7,1]. c). estimate the state's population in 2008

OpenStudy (a_clan):

a). P(x) = 4m *(1+0.033x) c). put x=8 in above equation to get population in 2008

OpenStudy (anonymous):

For exponential growth, use the formula, P = C (1 + r) ^t P = future value C = initial population r = growth rate (expressed as a fraction: eg. 0.06) t = number of years of growth Source: http://math2.org/math/general/interest.htm In 2004 (t=4), the population will be P = 4(1.033)^4 = about 4.55 million.

OpenStudy (anonymous):

thank u